OpenStudy (anonymous):

find the orthogonal projection of u = -i+j+k onto v= 2i+j-3k

OpenStudy (anonymous):

proj of u onto v = (u.v)/(v.v) * v u=(-1,1,1) and v=(2,1,-3) You can do it!

OpenStudy (anonymous):

lemme give it a go real quick

OpenStudy (anonymous):

is that v dot v or v^2?

OpenStudy (anonymous):

or ||v||^2?

OpenStudy (anonymous):

\[\frac{ u•v }{ v•v } v\]

OpenStudy (anonymous):

So yes, u dot v over v dot v

OpenStudy (anonymous):

so my final answer should be -4(-i+j+k) / 3 ?